# 0.5 eV QCD Axion Cosmology

### v48

@@ Abstract
A simple yet compelling physical picture is proposed for the nature of dark matter, dark energy, and the Big Bang. The proposal leads to a prediction with 0.53 parts per million precision for the 0.5 eV rest-mass energy of the quantum chromodynamic (QCD) axions that form the dark matter within this picture. A straightforward resolution of the Hubble tension follows within this picture from the agreement between the prediction for the present expansion rate of the universe and the best present determination from observations of the late universe.
@@ 1 Introduction
In 2022 the Hubble tension between late-universe observations of the present expansion rate (A. G. Riess et al., 2022) and predictions within the standard cold dark matter cosmology with cosmological constant ($\rm\Lambda CDM$}) based on early-universe observations (Planck Collaboration et al., 2020; S. Aiola et al., 2020) reached the conventional $5.0\sigma$ threshold for the discovery of a new phenomenon in astrophysics. Several years before this conventional threshold was reached, astrophysicists were already stimulated by the accumulating evidence for the existence of a significant tension, and they began suggesting innovative alternative cosmologies that might reconcile early- and late-universe observations (V. Poulin et al., 2019). In 2024 the tension was confirmed on the late-universe side with James Webb Space Telescope (JWST) observations that rejected unrecognized crowding of Cepheid variable star photometry as an explanation of the Hubble tension at $8\sigma$ confidence (A. G. Riess et al., 2024a), and that validated the Hubble Space Telescope (HST) distance measurements on which the tension rests (A. G. Riess et al., 2024b).
@@
On the early-universe side, the Dark Energy Spectroscopic Instrument (DESI) first-year observations of the peak in the galaxy-galaxy two-point correlation function confirmed the Hubble tension in 2025, provided the results are interpreted within the standard $\Lambda{\rm CDM}$ cosmology (A. G. Adame et al., 2025). Similarly, in 2023 an independent combination of ground-based and space-based observations of the cosmic microwave background (CMB) power spectrum of angular fluctuations confirmed the tension, again within the framework of $\Lambda{\rm CDM}$ (L. Balkenhol et al., 2023). In 2024 the first comprehensive book-length review concluded that the Hubble tension remains an unsolved mystery (A. G. Riess, 2024).
@@
The purpose of this Letter is to propose a simple yet compelling physical picture for the nature of dark matter, dark energy, and the Big Bang. However, the picture offers a straightforward resolution of the Hubble tension, and this resolution is, therefore, the primary astrophysical motivation for considering the picture, despite the fact that, historically, the picture emerged from quite different motivations and considerations (See Section 2). In brief, the picture leads to the following high-precision prediction from first-principles for the value of the present expansion rate of universe (See Section 3 for details):
$H_{0}=71.939~5~(76)~{\rm km\cdot sec^{-1}\cdot Mpc^{-1}}~[{\rm 0.11~per~mille}].$ (1)
@@
The predicted expansion rate in Eq. (1) roughly agrees, within the percent-level uncertainty of the comparison, with the best present late-universe determination using the relation between redshift and distance for Type Ia supernovae by the SH0ES collaboration (A. G. Riess et al., 2025)
$H_{0}({\rm SH0ES})=73.18~(88)~{\rm km\cdot sec^{-1}\cdot Mpc^{-1}}~[{\rm 12~per~mille}].$ (2)
@@
Correspondingly, the predicted expansion rate in Eq. (1) differs sharply, by $10\sigma$, from the best present $\Lambda{\rm CDM}$ prediction using early-universe observations (Planck Collaboration et al., 2019, 2020)
$H_{0}(\Lambda{\rm CDM})=67.9~(4)~{\rm km\cdot sec^{-1}\cdot Mpc^{-1}}~[{\rm 5.9~per~mille}].$ (3)
@@
Within the physical picture presented in this Letter, the straightforward resolution of the Hubble tension is simply that the late-universe observations are basically correct (on this point, see also D. Scolnic et al. (2025)), as are the $\Lambda{\rm CDM}$ predictions for the value of the cosmological constant and for the value of the present energy density of dark matter in the universe, but that the value of the present energy density of baryons is more than double the value predicted by $\Lambda{\rm CDM}$ (See Section 3 for the actual values of the corresponding cosmological parameters and for the comparison with $\Lambda{\rm CDM}$).
@@
In a spatially flat universe, the value of the present expansion rate, $H_{0}=100h~{\rm km~sec^{-1}~Mpc^{-1}}$, is linked to the values of the cosmological parameters of the cosmological constant, $\Lambda$, the dark matter energy density, $u_{A}$, the baryon energy density, $u_{B}$, and the photon energy density, $u_{\gamma}$, by the simple relation (P. J. E. Peebles, 1993, pg. 100)
$h^{2}=\Omega_{A}h^{2}+\Omega_{B}h^{2}+\Omega_{\gamma}h^{2}+\Omega_{\Lambda}h^{2},$ (4)
@@
where the cosmological parameter is $\Omega_{X}h^{2}=(8\pi/3)(Gu_{X}/c^{2})(h/H_{0})^{2}$ for energy density $u_{X}$, with $X=A,B,\gamma$, and it is $\Omega_{\Lambda}h^{2}=(c^{2}\Lambda/3)(h/H_{0})^{2}$ for the cosmological constant $\Lambda$. To resolve the Hubble tension, we propose a new cosmology, based on dark matter in the form of fundamental particles known as axions (S. Weinberg, 1978; F. Wilczek, 1978), that leads to predictions for the values of the cosmological parameters, $\Omega_{A}h^{2}$ and $\Omega_{B}h^{2}$, which have the same precision as the cosmological parameter for photons, $\Omega_{\gamma}h^{2}=2.473~(2)\times 10^{-5}~[{\rm 0.88~per~mille}]$, set by the precision of the best present estimate of the CMB temperature $T_{\gamma}=2.725~5~(6)~{\rm K}~[{\rm 0.22~per~mille}]$ (D. J. Fixsen, 2009). Together with the first-principles prediction, within this axion cosmology, for the value of $\Omega_{\Lambda}h^{2}$, to an astounding precision of 22 parts per million (ppm), the prediction for $H_{0}$ given in Eq. (1) is obtained using Eq. (4).
@@
The linchpin of the analysis leading to this straightforward resolution of the Hubble tension is the following remarkable, numerically exact relation for the ratio of the energy density of baryons to the energy density of axions in the present universe (See Section 4 for the derivation)
@@
where the small dimensionless quantity, $\epsilon=(4m_{p}/m_{\alpha}-1)Y_{P}\approx 0.002$, corrects for Big Bang nucleosynthesis of primordial helium with nucleus mass, $m_{\alpha}$, and primordial abundance by mass, $Y_{P}$, and the relativistic correction is $\delta\approx 0.06$ (See Appendix C), while the numerically exact constant, $x_{k}=2.821~439~372$ (P. J. E. Peebles, 1993, pg. 159), gives the kinetic energy, $\check{E}=x_{k}kT_{\pi}$, for the axions on the peak of the black-body distribution at the temperature of the Big Bang, $T_{\pi}=y_{T}\left(\chi_{\rm QCD}\right)^{1/4}/k$, and the numerically exact constant, $y_{T}=(5/84)^{1/4}=0.493~938~274$, relates this temperature to the topological susceptibility of the quantum chromodynamic (QCD) vacuum, $\chi_{\rm QCD}=(m_{A}c^{2})^{2}f_{A}^{2}$ (S. Weinberg, 1978; F. Wilczek, 1978), given by the axion rest-mass energy, $m_{A}c^{2}$, and the axion decay constant, $f_{A}$. The value of the corresponding ratio within $\Lambda{\rm CDM}$ cosmology is roughly half as large as the value given in Eq. (5) (Planck Collaboration et al., 2019, 2020). This sharp difference resolves the Hubble tension, as shown in Eq. (1), and has implications for a wide range of astrophysical phenomena, including the formation of large-scale structure during the cross-over from radiation-domination to matter-domination (See Section 5), the synthesis of the nuclei of the chemical elements with low atomic mass during the Big Bang (See Section 6), and the spectrum of initial fluctuations in energy density at the Big Bang itself (See Section 7).
@@
Finally, a wide range of astrophysical observations have been used over the past five decades to “constrain” or even “exclude” axion dark matter with the rest-mass energy given in Eq. (6) and the axion-photon coupling given in Eq. (7) (See Section 8). Most recently, in 2025, blank-sky observations with the JWST near-infrared spectrometer NIRSpec and mid-infrared spectrometer MIRI (R. Janish & E. Pinetti, 2025; S. Roy et al., 2025) were used to derive strong limits on the axion-photon coupling constant, $g_{A\gamma\gamma}$, in the mass range 0.1 to 4 eV, that “exclude,” with much greater than 95% confidence, these predicted values (E. Pinetti, 2025). Yet, more than a decade ago, in 2014, observations of horizontal branch stars in globular clusters, combined with stellar evolution that includes the core cooling effects from axion emission due to photon-to-axion conversion in the helium-burning cores of these stars, had already been used to derive a strong “upper bound” on the value of the axion-photon coupling, $g_{A\gamma\gamma}<0.66\times 10^{-10}~{\rm GeV}^{-1}$, independent of the value of $m_{A}$, which “excludes,” with just over 95% confidence, the predicted value in Eq. (7) (A. Ayala et al., 2014).

### v49

@@ Abstract
A simple yet compelling physical picture is proposed for the nature of dark matter, dark energy, and the Big Bang. The proposal leads to predictions, from first-principles with high precision, for the values of the Hubble constant, cosmological constant, and matter abundance in the universe. Early-universe observations yield a value for the redshift of matter-radiation equality, within the standard cold dark matter cosmology, that roughly matches the redshift where the equation of state crosses over from radiation-like to matter-like for the quantum chromodynamic (QCD) axion particles, with rest-mass energy per particle around 0.5 eV, and with number density six times that of the photons made in the Big Bang, that form the dark matter and that dominate the early-universe expansion dynamics within the proposed cosmology. Late-universe observations suggest a value for the Hubble constant that agrees, within the percent-level uncertainty of the comparison, with the value predicted by 0.5 eV QCD axion cosmology. Observations near cosmic noon, in the middle-universe, show pleasing agreement with the predicted values for the cosmological constant and matter abundance. Bolstered by this broad range of observational support, we re-visit the conventional astrophysical assumptions that have been used to rule out, constrain, and exclude 0.5 eV QCD axion dark matter for the past half century.
@@ I Introduction
In 2022 the Hubble tension between late-universe observations of the present expansion rate (A. G. Riess et al., 2022) and predictions within the standard cold dark matter cosmology with cosmological constant ($\rm\Lambda CDM$}), based on early-universe observations (Planck Collaboration et al., 2020; S. Aiola et al., 2020), reached the conventional $5.0\sigma$ threshold for the discovery of a new phenomenon in astrophysics. Several years before this conventional threshold was reached, astrophysicists were already stimulated by the accumulating evidence for the existence of a significant tension, and they began suggesting innovative alternative cosmologies that might reconcile early- and late-universe observations (V. Poulin et al., 2019). In 2024 the tension was confirmed on the late-universe side with James Webb Space Telescope (JWST) observations that rejected unrecognized crowding of Cepheid variable star photometry as an explanation of the Hubble tension at $8\sigma$ confidence (A. G. Riess et al., 2024a), and that validated the Hubble Space Telescope (HST) distance measurements on which the tension rests (A. G. Riess et al., 2024b).
@@
On the early-universe side, the Dark Energy Spectroscopic Instrument (DESI) first-year observations of the peak in the galaxy-galaxy two-point correlation function confirmed the Hubble tension in 2025, provided the results are interpreted within the standard $\Lambda{\rm CDM}$ cosmology (A. G. Adame et al., 2025). Similarly, in 2023 an independent combination of ground-based and space-based observations of the cosmic microwave background (CMB) power spectrum of angular fluctuations confirmed the tension, again within the framework of $\Lambda{\rm CDM}$ (L. Balkenhol et al., 2023). In 2024 a comprehensive book-length review concluded that the Hubble tension remains an unsolved mystery (A. G. Riess, 2024).
@@
The purpose of this Letter is to propose a simple yet compelling physical picture for the nature of dark matter, dark energy, and the Big Bang (See Section VI). The picture offers a straightforward resolution of the Hubble tension through the following high-precision prediction from first-principles for the value of the present expansion rate of the universe:
$H_{0}=71.939~5~(76)~{\rm km\cdot s^{-1}\cdot Mpc^{-1}}~[{\rm 0.11~per~mille}].$ (1)
@@
The predicted expansion rate in Eq. (1) roughly agrees, within the percent-level uncertainty of the comparison, with the best present late-universe determination, by the SH0ES collaboration, based on the relation between redshift and distance, within a flat, isotropic, homogeneous, expanding-universe cosmology (P. J. E. Peebles, 1993, pg. 73), for Type Ia supernovae (SNe Ia), calibrated both by Cepheid variables and by the tip of the red giant branch (TRGB), using both JWST and HST observations (A. G. Riess et al., 2025):
$H_{0}({\rm SH0ES})=73.18~(88)~{\rm km\cdot s^{-1}\cdot Mpc^{-1}}~[{\rm 12~per~mille}].$ (2)
@@
Correspondingly, the predicted expansion rate in Eq. (1) differs sharply, by $10\sigma$, from the best present $\Lambda{\rm CDM}$ prediction, based on early-universe observations (Planck Collaboration et al., 2019, 2020):
$H_{0}(\Lambda{\rm CDM})=67.9~(4)~{\rm km\cdot s^{-1}\cdot Mpc^{-1}}~[{\rm 5.9~per~mille}].$ (3)
@@
Within the physical picture presented in this Letter, the straightforward resolution of the Hubble tension is simply that the late-universe observations are basically correct (on this point, see also D. Scolnic et al. (2025)), as are the $\Lambda{\rm CDM}$ predictions (Planck Collaboration et al., 2019, 2020) for the values of the cosmological parameters for the cosmological constant, $\Omega_{\Lambda}(\Lambda{\rm CDM})=0.692~(5)~[{\rm 5.2~per~mille}]$, and the present energy density of dark matter in the universe, $\Omega_{A}h^{2}(\Lambda{\rm CDM})=0.118~82~(86)~[\rm 7.2~per~mille]$, but that the baryon abundance, $\Omega_{B}h^{2}=0.047~095~(31)~[0.66~{\rm per~mille}]$ (See Section VII for derivation), is more than double the value predicted by $\Lambda{\rm CDM}$: $\Omega_{B}h^{2}(\Lambda{\rm CDM})=0.022~46~(13)~[5.8~{\rm per~mille}]$ (Planck Collaboration et al., 2019, 2020).
@@
Within the proposed cosmology, as in $\Lambda{\rm CDM}$, space grows with time in a manner described by a spatially flat, homogeneous, isotropic, expanding-universe solution to the field equations of general relativity with cosmological constant term (P. J. E. Peebles, 1993, pg. 62, 65, 70–76, 265–268). Just as in $\Lambda{\rm CDM}$, then, the present value of the growth rate, $H_{0}=100h~{\rm km~s^{-1}~Mpc^{-1}}$, is given by the values of the cosmological parameters for the cosmological constant, $\Lambda$, the energy density of cold dark matter, $u_{A}=m_{A}c^{2}n_{A}$, the baryon energy density, $u_{B}$, and the photon energy density, $u_{\gamma}$, using the simple relation (P. J. E. Peebles, 1993, pg. 100),
$h=\sqrt{\Omega_{A}h^{2}+\Omega_{B}h^{2}+\Omega_{\gamma}h^{2}+\Omega_{\Lambda}h^{2}},$ (4)
@@
where the cosmological parameter is $\Omega_{X}h^{2}=(8\pi/3)(Gu_{X}/c^{2})(h/H_{0})^{2}$ for energy density $u_{X}$, with $X=A,B,\gamma$, and it is $\Omega_{\Lambda}h^{2}=(c^{2}\Lambda/3)(h/H_{0})^{2}$ for the cosmological constant $\Lambda$. To resolve the Hubble tension, we propose a new cosmology, 0.5 eV QCD axion cosmology, based on dark matter in the form of fundamental particles, known as quantum chromodynamic (QCD) axions (S. Weinberg, 1978; F. Wilczek, 1978), having rest-mass energy per particle, $m_{A}c^{2}\approx 0.5~{\rm eV}$ (See Appendix B for derivation), and having number density, $n_{A}/n_{\gamma}=6$ (See Appendix A for derivation), that is six times the number density of photons in the CMB, $n_{\gamma}\approx 411~{\rm cm^{-3}}$ (P. J. E. Peebles, 1993, pg. 158–159).
The proposed 0.5 eV QCD axion cosmology leads to predictions, from first-principles with high-precision, for the values of the cosmological parameters, $\Omega_{A}h^{2}$ and $\Omega_{B}h^{2}$, which have similar precision as the cosmological parameter for photons, $\Omega_{\gamma}h^{2}=2.473~(2)\times 10^{-5}~[{\rm 0.88~per~mille}]$ (P. J. E. Peebles, 1993, pg. 100, 137), set by the precision of the best present estimate of the CMB temperature, $T_{\gamma}=2.725~5~(6)~{\rm K}~[{\rm 0.22~per~mille}]$ (D. J. Fixsen, 2009) (See Section VII for derivation):
$\Omega_{A}h^{2}$ $=$ $0.117~949~(78)~[0.66~{\rm per~mille}]$
$\Omega_{B}h^{2}$ $=$ $0.047~095~(31)~[0.66~{\rm per~mille}].$ (5)
The prediction within 0.5 eV QCD axion cosmology for $H_{0}$, given in Eq. (1), is obtained using Eq. (4), by combining the values for $\Omega_{A}h^{2}$ and $\Omega_{B}h^{2}$, given in Eqs. (I), with the first-principles, high-precision prediction for the value of $\Omega_{\Lambda}h^{2}$ (See Section VII and Appendix E for derivation):
$\Omega_{\Lambda}h^{2}=0.352~460~(76)~[\rm 22~{\rm ppm}].$ (6)
@@
The linchpin of the analysis leading to this straightforward resolution of the Hubble tension is the following remarkable, numerically exact relation for the ratio of the energy density of baryons to the energy density of axions in the present universe (See Section VIII for the derivation):
@@
where the small dimensionless quantity, $\epsilon=(4m_{p}/m_{\alpha}-1)Y_{P}\approx 0.002$ (See Appendix C), corrects for Big Bang nucleosynthesis of primordial helium with nucleus mass, $m_{\alpha}$, and with primordial abundance by mass, $Y_{P}$, and the relativistic correction is $\delta\approx 0.06$ (See Appendix C), while the numerically exact constant, $x_{k}=2.821~439~372$ (P. J. E. Peebles, 1993, pg. 159), gives the kinetic energy, $\check{E}=x_{k}kT_{\pi}$, for the axions on the peak of the black-body distribution at the temperature of the Big Bang, $T_{\pi}=y_{T}\left(\chi_{\rm QCD}\right)^{1/4}/k$, and the numerically exact constant, $y_{T}=(5/84)^{1/4}=0.493~938~274$ (See Appendix B), relates this temperature to the topological susceptibility of the quantum chromodynamic (QCD) vacuum, $\chi_{\rm QCD}=(m_{A}c^{2})^{2}f_{A}^{2}$ (S. Weinberg, 1978; F. Wilczek, 1978), given by the axion rest-mass energy, $m_{A}c^{2}$, and the axion decay constant, $f_{A}$. The value of the corresponding ratio within $\Lambda{\rm CDM}$ cosmology is roughly half as large as the value given in Eq. (7) (Planck Collaboration et al., 2019, 2020). This sharp difference resolves the Hubble tension, as shown in Eq. (1), and has implications for a wide range of astrophysical phenomena, including the formation of large-scale structure during the cross-over from radiation-domination to matter-domination (See Section II), the synthesis of the nuclei of the chemical elements with low atomic mass during the Big Bang (See Section III), and the spectrum of initial fluctuations in energy density at the Big Bang (See Section IV).
Figure 1: Distance modulus, $\mu(z)=5~{\rm mag}\times\log_{10}(D(z)/{\rm Mpc})+25~{\rm mag}$, grows with redshift, $z$, in 0.5 eV QCD axion cosmology (teal line, ) in a way that is predicted from first-principles with high-precision by Eq. (8) and that lands on top of observations of 1635 high-$z$ SNe Ia from the Dark Energy Survey (DES) and 194 low-$z$ SNe Ia from the CfA/CSP + Foundation sample (violet pluses, +) (M. Vincenzi et al., 2024; DES Collaboration et al., 2024).
Looking back through the late-universe toward “cosmic noon” near redshift $z\approx 1$ (Fig. 1), the growth rate of space with time, $H(z)$, within 0.5 eV QCD axion cosmology, drives the growth of the luminosity distance, $D(z)$ (P. J. E. Peebles, 1993, pg. 319, 328), in a way that agrees with the Dark Energy Survey (DES) five-year observations of 15001500+ high-$z$ Type Ia supernovae (SNe Ia) (M. Vincenzi et al., 2024; DES Collaboration et al., 2024):
$D(z)$ $=$ $(1+z)\int_{0}^{z}dz\frac{c}{H(z)}=c\frac{h}{H_{0}}(1+z)\int_{0}^{z}dz\frac{1}{\sqrt{\Omega_{\Lambda}h^{2}+\Omega_{M}h^{2}(1+z)^{3}}},$ (8)
where $\Omega_{M}=1-\Omega_{\Lambda}-\Omega_{\gamma}=\Omega_{A}+\Omega_{B}=0.318~92~(14)~[0.44~{\rm per~mille}]$, gives the matter abundance in the universe. Together with the resolution of the Hubble tension, the agreement with observations out to cosmic noon for the growth of space with time, shown in Fig. 1, gives a strong sign that 0.5 eV QCD axion cosmology well describes the expansion of the universe back, at least, to the time when the universe was around half of its present size. Successful comparison with astrophysical observations of primordial deuterium abundance in the absorption line spectroscopy of quasars (R. J. Cooke et al., 2018) at redshifts around $z\approx 2$ (See Fig. 3), and of core helium abundance during star formation in globular clusters (See Fig. 4), push the agreement with 0.5 eV QCD axion cosmology into the epoch of Big Bang nucleosynthesis, $z_{\rm BBN}\approx 3\times 10^{8}$, while agreement with early-universe observations (See Fig. 2) probing the redshift of “matter-radiation equality,” $z_{\rm eq}\approx 3\times 10^{3}$, completes the story of observational support for the proposed cosmology.
@@
Finally, a wide range of astrophysical observations have been used over the past five decades to “constrain” or even “exclude” axion dark matter with the rest-mass energy given in Eq. (9) and the axion-photon coupling given in Eq. (10) (See Section V). Most recently, in 2025, blank-sky observations with the JWST near-infrared spectrometer NIRSpec and mid-infrared spectrometer MIRI (R. Janish & E. Pinetti, 2025; S. Roy et al., 2025) were used to derive strong limits on the axion-photon coupling constant, $g_{A\gamma\gamma}$, in the mass range 0.1 to 4 eV, that “exclude,” with much greater than 95% confidence, these predicted values (E. Pinetti, 2025). Yet, more than a decade ago, in 2014, observations of horizontal branch stars in globular clusters, combined with stellar evolution that includes the core cooling effects from axion emission due to photon-to-axion conversion in the helium-burning cores of these stars, had already been used to derive a strong “upper bound” on the value of the axion-photon coupling, $g_{A\gamma\gamma}<0.66\times 10^{-10}~{\rm GeV}^{-1}$, independent of the value of $m_{A}$, which “excludes,” with just over 95% confidence, the predicted value in Eq. (10) (A. Ayala et al., 2014).